For the system below: a) solve the system by the linear combination method. Show all work carefully. B) check your answer. X-2y=3 x+y=2 (y+3). Please having a bit of trouble on this one!
Answer by Alan3354(69443)

X-2y=3 --> x = 2y+3
x+y=2 (y+3) --> x+y = 2y+6 --> x = y+6
-----------
x = 2y+3
x = y+6
-------
2y + 3 = y + 6
y + 3 = 6
y = 3
------------
x = 9